OpenStudy (anonymous):

find the simplification of a^2a^7

OpenStudy (anonymous):

Use the law of exponents: \[e^me^n = e^{m+n}\]

OpenStudy (anonymous):

\[\Large a^ma^n = a^{m + n}\] if m = 2 and n = 7, what is the result?

OpenStudy (anonymous):

i dont get what your asking we havent been introduced to this formula yet.

OpenStudy (anonymous):

a^2 = a * a a^7 = a * a * a * a * a * a * a So a^2 * a^7 = a * a * a * a * a * a * a * a * a = a^9

OpenStudy (anonymous):

ooh! thank you that helps soo much!